OpenStudy (angelab97):

if each of the four letters in the sum below represents a different digit, which of the following cannot be a value of A? AB + BA ------ CDC 1. 6 2. 5 3. 4 4. 3 5. 2 How do you get the answer to this easily? Can anyone explain it thoroughly?

OpenStudy (anonymous):

Try setting A to each of those numbers, where A is different from B,C, and D. You should find that only one A value is possible.

OpenStudy (angelab97):

thanks but i tried that and i wasn't very successful. My book says that voice 5 is correct but i don't understand

OpenStudy (anonymous):

You can also solve this with a system of equations I believe.

OpenStudy (anonymous):

Hmm. This problem is kinda tricky. Hold on while I think about this a minute.

OpenStudy (anonymous):

Well I misread the question for one thing XD. ONE of those options is what A CANNOT be. So from the right-most collumn: B + A = C + k The next collumn: [A + B + k) ]%10 = D where k is what's carried from the first collumn and %10 is the remainder when divided by 10.

OpenStudy (angelab97):

thank you so much

OpenStudy (anonymous):

Did you find an answer? I'm still trying to work through it.

OpenStudy (angelab97):

i think the easiest way is to plug in for the numbers. I started with the last choice and plugged 2 for A and 8 for B and see why 2 is the answer

OpenStudy (angelab97):

but my answer changed when i plugged in 4..that's why I'm confused

OpenStudy (anonymous):

Using your method: 2B +B2 ---- CDC I don't see why any A value is impossible...

OpenStudy (anonymous):

Bah, I refreshed the page and my long post was lost. Hold on one more second... XD

OpenStudy (anonymous):

I think C must be 1 because: 1) you're using a base-10 number system(when you add two things that make ten you carry to the next collumn) 2) adding two 2-digit numbers where A and B are different, the largest carry you can have in the 2nd collum is 1. Proof: Let A = 9, and B = 8 (Or vise versa) (8 and 9 are chosen because they are the biggest numbers in a base-10 number system) 89 + 98 = 187

OpenStudy (anonymous):

So in my equations above k can only equal 0 or 1 and C = 1. This still leaves A, B and D. with only two equations...

OpenStudy (angelab97):

got it!

OpenStudy (anonymous):

How? I'm perplexed by this question!

OpenStudy (angelab97):

i plugged in random numbers until i got my answer. Sadly, this long problems is a PSAT sample question! It would've been miserable if i spent all my time on this problem.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4e2fa36d0b8ba7b2da403dbf Looks like this has been asked before XD

OpenStudy (anonymous):

Looks like there is no easy way to answer this question without trying all of the possibilities. But don't be discouraged, I'm in third year university XD.

OpenStudy (angelab97):

Thanks a lot..i guess all of the choices give you 121 except the last one